OpenStudy (anonymous):

Verify that the trigonometric function is an identity. cos theta ---------- = cos theta csc theta I don't understand this stuff at all and can't verify it.

OpenStudy (jhannybean):

\[\frac{\cos(\theta)}{\csc(\theta)} = \cos(\theta)\] you know that csc(theta)= 1/sin(theta) yes?

OpenStudy (jhannybean):

So you want to evaluate one side of the equation to match the other side.

OpenStudy (loser66):

It is not an identity, how to verify??